# Neural timescales from a computational perspective

### v1

@@ Abstract
Timescales of neural activity are diverse across and within brain areas, and experimental observations suggest that neural timescales reflect information in dynamic environments. However, these observations do not specify how neural timescales are shaped, nor whether particular timescales are necessary for neural computations and brain function. Here, we take a complementary perspective and synthesize three directions where computational methods can distill the broad set of empirical observations into quantitative and testable theories: We review (i) how data analysis methods allow us to capture different timescales of neural dynamics across different recording modalities, (ii) how computational models provide a mechanistic explanation for the emergence of diverse timescales, and (iii) how task-optimized models in machine learning uncover the functional relevance of neural timescales. This integrative computational approach, combined with empirical findings, would provide a more holistic understanding of how neural timescales capture the relationship between brain structure, dynamics, and behavior.
@@ 1 Introduction
Ongoing neural activity unfolds across a wide range of timescales that are suggested to be related to the large diversity of behavioral timescales and those of the behavior-relevant inputs to the nervous system. These timescales are often characterized by the decay rate of a neural signal’s autocorrelation function. While precise definitions differ, recent experimental studies of neural timescales present a consistent picture across different recording modalities, species, and cognitive tasks: timescales increase along the cortical hierarchy, are diverse within each brain area, and are often correlated with behavioral variables during cognitive tasks (for recent reviews of empirical evidence, see[1, 2, 3]). As such, this fundamental quantity—the characteristic time constant of neuronal dynamics—may be an important signature of neural circuits across the brain. Importantly, neural timescales are suggested to take on a “representational” role, i.e., the diversity of timescales across the brain mirrors the multiple fluctuation timescales of relevant information in a dynamic environment, which is leveraged by the brain for perception, memory, planning, and action. However, while the large body of experimental observations demonstrates the relevance of neural timescales, they often blur crucial differences between signal modalities, and do not specify the mechanisms underlying how neural timescales arise, nor whether they are necessary (or sufficient) for representation and transformation of information in the brain.
@@
Here, we take a computational perspective and review how models of the brain—when constrained by multi-modal observations of neural timescales—can help provide a mechanistic understanding of neural timescales complementary to experimental investigations. Specifically, we discuss three research directions for how computational methods can distill the broad set of empirical observations into quantitative and testable theories, formulated as the following questions (Fig. 1):
@@
  • “Neural timescales” are measured from different physiological signals (e.g., spikes, field potentials, calcium imaging, fMRI BOLD) and quantified using different methods: how do these measurements relate and differ from one another? (Section 2)
@@
  • Which cellular and circuit processes contribute to the diversity of observed neural timescales, and how can mechanistic models of brain dynamics help propose and exclude candidate mechanisms? (Section 3)
@@
  • When are “representational neural timescales” that mirror environmental timescales necessary (or sufficient) for supporting task-relevant computations, and how to use functional models from machine learning (in particular, recurrent neural networks) to explore such hypotheses? (Section 4)
We advocate using an integrative perspective, combining computational and empirical approaches, to understand how neural timescales provide a link between brain structure, dynamics, and behavior and discuss studies that laid stepping stones to build such a perspective.
@@
Fig. 1: Graphical summary. This review discusses three main directions, demonstrating conceptually: (i) how estimation methods contribute to advancing measuring timescales from different modalities, (ii) how mechanistic models lead to uncovering the underlying circuit origins of timescales, and (iii) how functional modeling facilitate investigating their computational relevance.
@@ 2 Estimating timescales from neural recordings
Neural timescales are estimated from different recording modalities, such as spiking activity[4, 5, 6, 7, 8], LFP[9], ECoG[10, 11], calcium imaging[12, 13], fMRI BOLD[14, 15, 16], using a variety of estimation methods[4, 11, 7, 17]. Different modalities capture neural activity with distinct spatiotemporal resolutions (e.g., fast and local electrophysiological recordings versus slow and global fMRI BOLD), reflect different underlying physiological mechanisms (e.g., neuronal membrane potential versus metabolic processes), and are described by different mathematical models (e.g., point processes versus continuous signals). At the same time, estimation methods for corresponding timescales rely on distinct mathematical assumptions that may affect the subsequent results. While studies of neural timescales often concentrate on the similarity of findings across different modalities and methods, both the recording modality and the analysis method can introduce constraints and biases on estimated timescales that need to be carefully considered when assessing the generalisability of results.
Fig. 2: Estimating neural timescales from different data modalities using various methods. a. Timescales estimated from neural activity measured by different recording modalities increase along the visual cortical hierarchy, but their values differ by orders of magnitude. MT - medial-temporal area, LIP - lateral intraparietal area, OFC - orbitofrontal cortex, LPFC - lateral prefrontal cortex, ACC - anterior cingulate cortex (data from [4, 11, 16]). b. Different methods for estimating neural activity timescales ($\tau$) generally vary across two axes: (i) Model explicitness and complexity (vertical), (ii) Number of features they capture in data (horizontal): from the approximate decay rate $\tau$, to capturing autocorrelation shape by including a baseline representing slow processes across trials or oscillatory components, to including task-related components and prior knowledge about the data (panels are adapted from [18, 4, 19, 7, 17]).
2.1 The impact of recording modality on estimated timescales
@@
One of the most robust observations across various recording modalities is the hierarchical organization of neural timescales, i.e., timescales being faster in primary sensory and motor areas, and slower in transmodal association areas. This hierarchy has been observed in spiking activity[4, 7], LFP[9], ECoG[11, 10], calcium imaging[12, 13] and fMRI[14, 15, 20, 16]. However, despite the similar hierarchical organization, these observations have clear distinctions (Fig. 2a).
@@
Timescales measured from different modalities are affected by the biophysical properties of recorded signals and physical constraints in the recording setup. Spiking activity timescales are often in the range of tens to hundreds of milliseconds[4, 7], whereas fMRI BOLD timescales are in the range of a few seconds[14, 15, 16]. For the same stimulus, the hemodynamic evoked response is generally slower than spiking activity[21, 22]. Moreover, the currently used slow fMRI scan rates may further limit the detection of faster timescales in neural activity[23, 24]. Similar issues can affect timescales measured from calcium imaging: The time constant of calcium release, as well as the time constant of the calcium indicator and the imaging scan rate, can cause the same underlying neural dynamics to appear as having a different timescale[25]. Thus, when interpreting their potential functional relevance, it is important to distinguish whether estimated timescale differences arise from underlying population dynamics, biophysics of the measurement process, or both.
On the other hand, neural timescales estimated from different recording modalities may reflect distinct underlying neurophysiological processes and computations that are of further interest. fMRI BOLD is often associated with population activity [21], hence, its timescales may reflect population spiking timescales. Estimated timescales from LFP[9] and ECoG[11]; however, are approximately ten times faster than spiking activity in similar brain regions, which seems to be at odds with the notion that LFP and ECoG signals are slower aggregate signals than spiking activity[21, 26]. This observation suggests that timescales estimated from different modalities may arise from distinct but potentially correlated physiological processes, e.g., fast transmembrane and synaptic currents versus slow recurrent interactions[8]. In fact, when explicitly accounting for multiple timescales in the estimation method[17], both fast and slow timescales can be uncovered from spiking activity[8]. Therefore, estimation methods that consider multiple timescales can potentially incorporate modality- or even behavior-specific time constants explicitly (as we discuss in the next section[17, 7]).
The observed hierarchical organization of intrinsic timescales presents differences across studies, potentially stemming from differences across species and methodologies. Unlike primate studies[4, 11, 16], studies on the hierarchy of timescales in mouse cortex have reported inconsistent results. Intrinsic timescales measured from calcium imaging[12] and fMRI[20] display a hierarchical structure from visual to prefrontal areas. However, in spiking activity, such a hierarchy was initially observed only in timescales of stimulus-evoked activity [27, 28] and not in spontaneous activity [27], though recently it was shown that by analyzing longer periods of recorded neural activity, the hierarchical structure can also be recovered in spontaneous activity[29]. These results suggest that multiple timescales might be present in spontaneous activity, but only some follow the anatomical hierarchy. Moreover, even in primate studies, the hierarchical ordering of brain areas based on timescales does not exactly match across different modalities[4, 11, 16] (Fig. 2a). Studying such subtle but consistent differences may help to better understand the inherent biophysical differences between various recording modalities, and cross-species differences in cortical architecture.
@@ 2.2 Different methods for estimating timescales
Methods for estimating timescales from neural activity usually quantify the decay rate of the signal’s autocorrelation function. However, each method relies on different theoretical assumptions and incorporates different features of neural dynamics (Fig. 2b). For instance, they could explicitly include oscillatory components that are often present in neural activity and can be modulated by task variables, which can alter the autocorrelation shape. Here, we catalog existing methods along two axes: (i) the explicitness of their fitted model and their computational complexity, and (ii) the number of features they capture in data (Fig. 2b).
@@
The first group of methods does not assume any strict mathematical form for the autocorrelation shape. These methods characterize the decay rate as, for example, the time lag where autocorrelation drops to $1/e$[30], the (half of) full width at half maximum[14], or the integral of positive autocorrelation values[18, 16]. We thus refer to these methods as “model-free” methods. Computing such measures is straightforward, and they are typically estimated directly from the data. However, such model-free methods do not capture the full shape of data autocorrelation. As a result, in the presence of complex temporal features in neural activity beyond a single exponential decay, different methods can give different results on the same data (for example, see [30, 31]).
@@
The second group of methods assumes a mathematical function for the shape of neural data autocorrelation to capture the relevant data features and estimate the timescale from the parameters of the fitted function (i.e., model-based). The simplest assumption is that the autocorrelation of neural activity follows an exponential decay function, $\textrm{AC}(t)=e^{-t/\tau}$, where decay time constant $\tau$ reflects the timescale of underlying neural dynamics. This is equivalent to the assumption of obtaining a Lorentzian function as a power spectral density in the frequency domain, i.e., $\textrm{PSD}(f)\sim(f^{2}+f_{k}^{2})^{-1}$, where the knee frequency $f_{k}$ is translated to timescale as $\tau=(2\pi f_{k})^{-1}$ [11]. The exponential decay (Lorentzian function) can be further augmented with additional features to better capture the complexity of temporal dynamics. For example, the autocorrelation of neural dynamics can exhibit more than one timescale[4, 32, 7, 8]. Very slow timescales can also arise from firing rate fluctuations across different trials. Such a slow timescale can be accounted for by fitting the autocorrelation with an exponential decay function and an additional offset term corresponding to longer-term processes (e.g.,[4, 5], see[33] for a toolbox). Alternatively, multiple timescales can be estimated by fitting the PSD with a mixture of Lorenzian functions[32].
Neural activity can additionally contain oscillatory dynamics that appear as periodic fluctuations in the autocorrelation or peaks in the PSD, which can impact the estimated timescale or knee frequency, respectively. To overcome this problem, methods such as Spectral Parameterization (or, “FOOOF”) remove the oscillatory components from the PSD before estimating the knee frequency[34, 11, 35]. For example, oscillatory peaks can be modeled as a mixture of Gaussians and removed. Then, the timescale is estimated by fitting an extended form of the Lorentzian function as $\textrm{PSD}(f)=C(f^{\chi}+f_{k}^{\chi})^{-1}$, with a variable 1/f power-law exponent, allowing a more flexible fitting of the neural power spectra.
@@
Finally, a more accurate estimation of timescales can be obtained using a generative process capturing the statistics of neural time series[17, 7]. Such methods are particularly powerful since they can explicitly incorporate modality-specific properties and task variables. For instance, including task variables in a multi-timescale autoregressive process allows for disentangling intrinsic and task-induced timescales such as fluctuations induced by reward[7, 12, 28]. Moreover, generative models can capture experimental limitations, such as short trial durations that can lead to systematic underestimation of timescales[17].
@@
The parameters of the generative process can be directly fitted to neural time series[7] (e.g., via least squares or maximum likelihood) or can be used to instantiate a generative model in Bayesian inference methods[17, 36, 37, 38, 39, 40]. Bayesian methods aim to uncover the whole set of parameters that is compatible with the data (i.e., the posterior distribution) instead of obtaining the single best-fitting value of timescales (or other parameters), allowing them to better deal with the stochasticity of neural activity. For simple models, the posterior distribution can be computed analytically[36], but for more complex models, simulation-based methods such as Approximate Bayesian Computation (e.g., abcTau[17]) and simulation-based inference (SBI[38, 39, 41, 40]) can be applied to obtain approximate posterior distributions. In such methods, synthetic time series are produced from an explicit generative (statistical or mechanistic) model mimicking similar statistics as neural recordings. Then, by matching the summary statistics (e.g., autocorrelation, PSD, or time series) of synthetic data to the experimental data, we can estimate timescales using the parameters of the matched generative model.
@@
The generative model and summary statistics can be chosen flexibly, allowing them to capture different data features. It is possible to fit different generative models (e.g., models with different numbers of timescales or oscillatory components) and use Bayesian model comparison techniques to find the best models describing the neural data[17, 42]. Since the generative model can also incorporate modality-specific properties, such as timescale ranges, stochastic elements (e.g., spiking activity as a point process versus LFP as a continuous signal [43, 44, 45]), or even particular biophysical properties, it can be used to uncover possible models describing each modality. Therefore, these methods can incorporate various mechanistic assumptions and data features when estimating timescales, but this comes with additional computational costs, both in running the simulations and during posterior inference.

### v2

@@ Abstract
Neural activity fluctuates over a wide range of timescales within and across brain areas. Experimental observations suggest that diverse neural timescales reflect information in dynamic environments. However, how timescales are defined and measured from brain recordings vary across the literature. Moreover, these observations do not specify the mechanisms underlying timescale variations, nor whether specific timescales are necessary for neural computation and brain function. Here, we synthesize three directions where computational approaches can distill the broad set of empirical observations into quantitative and testable theories: We review (i) how different data analysis methods quantify timescales across distinct behavioral states and recording modalities, (ii) how biophysical models provide mechanistic explanations for the emergence of diverse timescales, and (iii) how task-performing networks and machine learning models uncover the functional relevance of neural timescales. This integrative computational perspective thus complements experimental investigations, providing a holistic view on how neural timescales reflect the relationship between brain structure, dynamics, and behavior.
@@ 1 Introduction
Ongoing neural activity unfolds across a wide range of timescales. These timescales are often characterized by the decay rate of a neural signal’s autocorrelation function. While precise definitions differ, recent experimental studies of neural timescales present a consistent picture across different recording modalities, species, and cognitive tasks: timescales increase along the cortical hierarchy, are diverse within each brain area, and are often correlated with behavioral variables during cognitive tasks (for recent reviews of empirical evidence, see [1, 2, 3]). As such, this fundamental quantity—the characteristic time constant of neuronal dynamics—may be an important signature of neural circuits across the brain. Importantly, neural timescales have been suggested to take on a “representational” role, i.e., the diversity of timescales across the brain mirrors the multiple fluctuation timescales of relevant information in a dynamic environment, which is leveraged by the brain for perception, memory, planning, and action [1, 2, 3]. However, while a large body of experimental observations demonstrates the relevance of neural timescales, crucial differences exist between how these timescales are defined and measured. Furthermore, it is experimentally challenging to probe the mechanisms underlying how neural timescales arise, and whether they are necessary (or sufficient) for representation and transformation of information in the brain.
@@
Supplementing experimental investigations, we here contribute a complementary perspective and review how computational methods have shed light on the origins and functions of neural timescales, while clarifying definitions for different types of timescales in the literature (Box 1). Specifically, we highlight three directions for how computational methods can disentangle different types of timescales and distill the broad set of empirical observations into quantitative and testable theories, formulated as the following questions (Fig. 1):
@@
  • “Neural timescales” are measured from different brain signals, during different tasks, and quantified using different methods: how do these measurements relate and differ from one another? (Section 2)
@@
  • Which cellular and circuit processes contribute to shaping measured timescales, and how can mechanistic models of brain dynamics help propose and exclude candidate mechanisms? (Section 3)
@@
  • When are “representational neural timescales” that mirror environmental timescales necessary (or sufficient) for supporting task-relevant computations, and how to use functional models from machine learning to explore such hypotheses? (Section 4)
Moreover, we review how recent advances in machine learning (ML) and artificial intelligence (AI) have opened new avenues to develop novel data analysis methods and computational models, getting us closer to an understanding of how neural timescales capture the relationship between brain structure, dynamics, and behavior.
[!ht] Box 1. Glossary: Different usages of “timescales” The definition of timescale varies in the literature across studies. Here, we briefly outline different notions of timescale that are commonly (and sometimes implicitly) used, with more detailed discussions throughout the Review:
Biophysical time constants:
Timescales can be thought to correspond to the passive membrane or adaptation time constants of single neurons. Similarly, timescales measured from field potential data may relate to time constants of synaptic or ion channel currents (Box 2).
Activity timescales in linear(ized) networks:
Another common notion refers to the decay time constant of population activity. Specifically, the output of a linear time-invariant system (i.e., a linear recurrent neural network) can be decomposed into independent modes of (complex) exponentials with different decay time constants (Box 3). This notion is often leveraged to interpret estimated timescales with the implicit assumption that network dynamics is (locally) linear, since “time constants” in nonlinear networks are state-dependent and thus ill-defined as a global, fixed property.
Response timescales:
Aside from the above “intrinsic” properties of the neural system under observation, measured timescales can also reflect timescales embedded in the input or be modulated by task and behavioral states. Response timescales could thus correspond to multiple timescales in a streaming sensory stimulus (e.g., a movie), or more generally, to upstream neural activity onto the downstream, recorded region (e.g., V1 to V4). They can also arise from modulations by top-down inputs during a behavioral task.
Any quantification of timescales from neural recordings is likely to reflect, to different degrees, a mixture of the above. Therefore, a key challenge moving forward for both computational methods and experimental design lies in separating those properties of neural systems from measurements of neural activity (Section 2), as well as understanding their roles in neural circuit dynamics and computation (Sections 3 and 4).
@@
Fig. 1: Graphical summary. This review discusses three main directions, demonstrating conceptually: (i) how estimation methods contribute to advancing the measurement of timescales from different experimental paradigms (e.g., recording modalities), (ii) how mechanistic models can help in uncovering the underlying circuit origins of different timescales, and (iii) how functional modeling facilitates investigation of their computational relevance.
@@ 2 Estimating timescales from neural recordings
Neural timescales have been estimated from data collected across various experimental paradigms, including different species, tasks, and recording modalities, as well as using different estimation methods (Table 1). While studies of neural timescales often concentrate on the similarity of their findings across different paradigms and methods, aforementioned differences also introduce assumptions and biases on estimated timescales that may lead to important conceptual deviations. In the following, we discuss these assumptions and biases and how they should be incorporated when assessing the generalizability of results.
Fig. 2: Diversity of neural timescales across and within brain areas. a. Average estimated timescales from spontaneous neural activity measured by different recording modalities consistently increase along the cortical hierarchy, but their values differ by orders of magnitude (data from [4, 5, 6]). b. Neural timescales are also diverse across individual neurons within one brain area (data from [7]). MT: medial-temporal area, LIP: lateral intraparietal area, OFC: orbitofrontal cortex, LPFC: lateral prefrontal cortex, DLPFC: dorsolateral prefrontal cortex, ACC: anterior cingulate cortex.
2.1 The impact of experimental paradigm on estimated timescales
@@
One of the most robust observations across different studies is the hierarchical organization of neural timescales, i.e., timescales being faster in primary sensory and motor areas, and slower in transmodal association areas. The hierarchy has been reported across various species (rodents and primates), recording modalities (spiking activity, calcium imaging, LFP, ECoG, and fMRI), and cognitive tasks (Table 1), suggesting that shared principles may give rise to these observations. However, there are also important distinctions that may be useful for uncovering the underlying mechanisms.
Modality Method Model-free Variants of exponential fit Autoregressive models Bayesian methods
Spiking activity Mendoza-Halliday et al. [8],  Wang et al. [9],  Meirhaeghe et al. [10],  Tang et al. [11] Zeisler et al. [12]∗∗,  Murray et al. [4],  Wasmuht et al. [13],  Cavanagh et al. [7],  Fascianelli et al. [14], Cirillo et al. [15],  Bernacchia et al. [16],  Fontanier et al. [17]∗∗,  Zeisler et al. [12]∗∗,  Siegle et al. [18]∗∗,  Rudelt et al. [19]∗∗,  Rudelt et al. [20],  Zeisler et al. [12]∗∗ Spitmaan et al. [21]∗∗,  Song et al. [22]∗∗,  Song et al. [22]∗∗ Zeraati et al. [23],  Zeraati et al. [24]∗∗,  Neophytou et al. [25],  Gonçalves et al. [26],  Gao et al. [27],  Vinogradov et al. [28]
Calcium imaging Çatal et al. [29]∗∗ Runyan et al. [30],  Imani et al. [31],  Piasini et al. [32]∗∗ Pinto et al. [33]∗∗ Nair et al. [34], Vinograd et al. [35],  Vinogradov et al. [28]
LFP, ECoG, EEG Honey et al. [36] Donoghue et al. [37],  Brake et al. [38], Chaudhuri et al. [39],  Gao et al. [5]∗∗,  Çatal et al. [29]∗∗, Cusinato et al. [40],  Gao et al. [5],  Manea et al. [41]∗∗
fMRI Hasson et al. [42],  Raut et al. [43],  Fallon et al. [44],  Watanabe et al. [45],  Shafiei et al. [46], Manea et al. [6],  Sethi et al. [47] Fallon et al. [44]
Table 1: Overview of timescale measurements in experimental literature. Timescales are measured using various recording modalities, estimation methods, and species (indicated by colors as human, non-human primate and rodent). * indicates studies that measured timescales from task-driven activity and ** from both spontaneous and task-driven activity. The rest measured timescales from spontaneous activity. Repeated entries indicate multiple methods, modalities, or species.
@@
Timescales estimated from different recording modalities increase along the cortical hierarchy, though their absolute values differ by orders of magnitude (Fig. 2a). These differences may stem from either biophysical properties of the circuits (e.g., time constants), or constraints in the experimental recording setup, such as sampling rate. Spiking activity timescales are often reported to be in the range of tens to hundreds of milliseconds [4, 21], whereas fMRI BOLD timescales are in the range of a few seconds [43, 44, 6]. Since BOLD signals partially reflect population spiking activity [48, 49] by measuring a lagged hemodynamic response [48], timescales estimated from fMRI may correspond to “slowed” population spiking timescales. Estimated timescales from LFP [41] and ECoG [5], on the other hand, are approximately ten times faster than spiking activity in similar brain regions, which seems to be at odds with the notion that LFP and ECoG are slower aggregate signals [48, 50]. This observation suggests that timescales estimated from different modalities may be dominated by distinct but potentially correlated physiological processes, e.g., fast transmembrane and synaptic currents versus slow recurrent network activity. In fact, when explicitly accounting for multiple timescales in the estimation method [23], both fast and slow timescales can be uncovered from spiking activity [24]. At the same time, currently used slow fMRI scan rates may further limit the detection of faster timescales in neural activity [51, 52]. Similar issues can affect timescales measured from calcium imaging: The time constant of calcium release, as well as the time constant of the calcium indicator and the imaging scan rate, can cause the same underlying neural dynamics to appear as having a different timescale [53]. Thus, when interpreting their potential functional relevance, it is important to distinguish whether estimated timescale differences arise from underlying neurophysiological properties, measurement process constraints, or both.
Neural timescales can also be estimated from different behavioral states and task conditions, resulting in two conceptually different quantities: (i) intrinsic timescales reflecting spontaneous fluctuations of neural activity (e.g., during fixation or rest), and (ii) response timescales arising from stimulus- or task-driven activity fluctuations. While both intrinsic and response timescales increase along the cortical hierarchy [4, 18, 20, 32, 6, 5, 41, 43, 21, 22, 54, 36, 42], they are not correlated within single neurons [21]. The reason behind this observation may be two-fold: First, single-neuron activity timescales are highly diverse within individual brain areas [7, 13, 14, 15], and even areas at the top of the hierarchy have many neurons with fast timescales (Fig. 2b). Various studies have shown that neurons with fast and slow timescales engage in different task epochs and are relevant to distinct types of computation (e.g., [7, 13, 14, 17], for a review see [1]). Second, neural timescales are state- and input-dependent, and change according to task demands [5, 24, 41, 17] or input timescales [32]. Thus, distinct mechanisms may underlie intrinsic and response timescales. In the following sections, we discuss how generative model-based methods can incorporate modality- and task-specific components when estimating neural timescales and how mechanistic models allow us to disentangle their underlying mechanisms, e.g., state-dependent timescales arising from nonlinear brain dynamics.
@@ 2.2 Different methods for estimating timescales
Extracting timescales from neural activity usually requires two steps: (i) computing a form of summary statistic that quantifies temporal fluctuations, and (ii) estimating timescales from the summary statistic (Fig. 3). For the summary statistic, most methods use some variant of time-lagged metrics, such as autocorrelation (AC, or its frequency domain equivalent, power spectral density, PSD) or delayed mutual information (MI, Fig. 3a, top). For example, intrinsic timescales are characterized by the decay rate of the spontaneous activity autocorrelation computed from individual experimental trials [4, 7, 13, 14, 16, 17, 18, 19, 24, 23, 5, 31, 32, 37, 39, 29, 6, 41, 44, 43, 45], whereas response timescales are quantified by the autocorrelation of trial-averaged activity post-stimulus [32, 18]. Response timescales can also be estimated with decoding methods that assess how long the encoding of a given stimulus or task event persists over time [30, 20, 19, 32]. Here, the summary statistic is often constructed using metrics such as decoding accuracy or predictability as a function of time (Fig. 3a, bottom) [30, 20, 19].
Fig. 3: Different methods for estimating neural timescales. a. Neural activity fluctuations can be characterized by various summary statistics such as the autocorrelation or delayed mutual information (AC / MI, top), or information content in neural activity about a preceding stimulus (bottom) (adapted from [30, 20, 19]). b. Methods for estimating timescales ($\tau$) from summary statistics generally vary along two axes: (i) Computational complexity (vertical), (ii) Number of features they capture in data (horizontal): from the approximate decay rate to capturing autocorrelation shape by including a baseline representing slow processes across trials or oscillatory components, to including task-related components and prior knowledge about the data (panels are adapted from [45, 4, 55, 21, 23]).
Using a computed summary statistic (e.g., AC or MI) as input, methods for estimating timescales rely on distinct theoretical assumptions and incorporate different features of neural dynamics. Here, we catalog existing methods along two axes: (i) computational complexity (often as a result of using more detailed models) and (ii) the number of features they capture in data (Fig. 3b). Note that, while in the following we mainly focus on quantifying timescales from the autocorrelation function, similar analyses can be applied to information-theoretic summary statistics.
@@
The first group of methods does not assume any strict mathematical form for the autocorrelation shape. These methods characterize the decay rate as, for example, the time lag where autocorrelation drops to $1/e$ [56] or 0.5 (i.e., half width at half maximum [43]), or the integral of positive autocorrelation values [45, 6]. We thus refer to these methods as “model-free”. Computing such measures is straightforward, and they are typically estimated directly from the data. However, such model-free methods do not consider the full shape of data autocorrelation. As a result, in the presence of complex temporal features in neural activity beyond a single exponential decay, different methods can give different results on the same data (for example, see [56, 57]).
@@
The second group of methods assumes a mathematical function for the autocorrelation to capture the relevant data features (i.e., model-based), and estimate timescales from the parameters of the fitted function (e.g., via least squares). The simplest assumption is that the autocorrelation of neural activity follows an exponential decay function, $\textrm{AC}(t)=e^{-t/\tau}$, where decay time constant $\tau$ reflects the timescale of underlying neural dynamics. This is equivalent to the assumption of obtaining a Lorentzian function as a power spectral density in the frequency domain, i.e., $\textrm{PSD}(f)\sim(f^{2}+f_{k}^{2})^{-1}$, where the knee frequency $f_{k}$ is translated to timescale as $\tau=(2\pi f_{k})^{-1}$ [5]. The exponential decay (or Lorentzian function) can be further augmented with additional features to better capture the complexity of temporal dynamics. For example, the autocorrelation of neural dynamics can exhibit multiple timescales [39, 21, 24] that may correspond to distinct neural mechanisms. In such cases, timescales can be estimated by fitting a linear mixture of exponential (or Lorentzian) functions [39, 19]. Moreover, very slow timescales can also arise from firing rate fluctuations across different trials, which can be accounted for by fitting the autocorrelation with an exponential decay function and an additional offset term corresponding to longer-term processes (e.g., [4, 7], see [58] for a toolbox). Neural activity can additionally contain oscillatory dynamics that appear as periodic fluctuations in the autocorrelation or peaks in the PSD, which can impact the estimated timescale or knee frequency, respectively. To overcome this problem, methods such as Spectral Parameterization [37, 5, 38] model oscillatory peaks as a mixture of Gaussians, which are removed before fitting an extended form of the Lorentzian function as $\textrm{PSD}(f)=C(f^{\chi}+f_{k}^{\chi})^{-1}$ (i.e., with a variable 1/f power-law exponent) to allow more flexible fitting of the neural power spectra.
@@
Finally, a more accurate estimation of timescales can be obtained using a generative process capturing the statistics of neural time series [23, 21, 25, 59]. Such methods are particularly powerful since they can explicitly incorporate modality-specific properties and task variables. For instance, including task variables in an autoregressive generative process has allowed for disentangling intrinsic and response timescales such as fluctuations induced by reward [21, 33, 22]. Moreover, multiple timescales can be explicitly modeled in the generative process, for instance, as a mixture of autoregressive processes with distinct timescales [23, 24, 21]. Generative models can also capture experimental limitations, such as short trial durations that affect the shape of data autocorrelation and can lead to systematic underestimation of timescales by model-free and model-based methods discussed earlier [23].
@@
The parameters of the generative process can be directly fitted to neural time series [21] (e.g., via least squares or maximum likelihood) or can be used to instantiate a generative model in Bayesian inference methods [23, 59, 25, 26, 60, 61, 27, 62]. Bayesian methods aim to uncover the whole set of parameters that is compatible with the data (i.e., the posterior distribution) instead of obtaining the single best-fitting value of timescales (or other parameters), allowing them to better deal with the stochasticity of neural activity. For simple models, the posterior distribution can be computed analytically [59], but for more complex models, simulation-based methods such as approximate Bayesian computation (e.g., abcTau [23]) and simulation-based inference (SBI [26, 60, 61, 63, 27]) can be applied to obtain approximate posterior distributions. In such methods, synthetic time series are produced from an explicit (statistical or mechanistic) generative model mimicking similar statistics as neural recordings. Then, by matching the summary statistics (e.g., autocorrelation, PSD, or time series) of synthetic data to the experimental data, we can estimate timescales using the parameters of the matched generative model.
@@
The generative model and summary statistics can be chosen flexibly, allowing them to capture different data features. It is possible to fit different generative models (e.g., models with different numbers of timescales or oscillatory components) and use Bayesian model comparison techniques to find the best models describing the neural data [23, 64]. Since the generative model can also incorporate modality-specific properties, such as timescale ranges, stochastic elements (e.g., spiking activity as a point process versus LFP as a continuous signal [65, 66, 67]), or even particular biophysical properties, it can be used to uncover possible models describing each modality. Therefore, while these methods often come with additional computational costs from performing simulations and posterior inference, they can incorporate important mechanistic assumptions and data features—both based on neurobiology and biophysics of the signal-generating process—when estimating timescales.